joshiscool:

Rohan is checking in poodles for a dog show. A miniature poodle must be between 10 in. and 15 in. at the shoulder. The body length must be no more than 16 in. In addition, the poodle’s shoulder height must be no more than 1 in. longer than its body length. The graph shows the feasible region, where x represents the poodle’s body length and y represents the shoulder height. Which ordered pairs meet all the constraints for a successful poodle measurement and make sense in context of the situation? Select each correct answer. (11, 9) (12, 10.5) (15, 11) (10, 12) (11, 14)

joshiscool:

@JustSaiyan

joshiscool:

@Elsa213

JustSaiyan:

@Vocaloid @Hero

joshiscool:

@Allison

Allison:

Where's the graph, @joshiscool ?

joshiscool:

how on

Allison:

Huh?

joshiscool:

hold on

joshiscool:

http://prntscr.com/gt7mbt

joshiscool:

@Allison

joshiscool:

joshiscool:

joshiscool:

Allison:

I think just B and C..I'm really not sure though. B and C are the only points that go inside the triangle. @Vocaloid

joshiscool:

Thank you

Allison:

What if I'm wrong?! D:

joshiscool:

we shall see

Allison:

*Starts sweating from pressure* xD

joshiscool:

What is the point slope form of the line with slope −14 that passes through the point (−2, 9) ? y+9=−14(x−2) y+2=−14(x−9) y−9=−14(x+2) y−2=−14(x+9)

joshiscool:

^

Allison:

Slope e.e I hate slope. Ermm @ThisGirlPretty @Falconmaster @Elsa213 @sammixboo @AshtynThatDrummerGirl @katherine @RawlsK29

RawlsK29:

let me look at this real quick

Allison:

Okay she got this from here~

RawlsK29:

Umm.. I think it is either B or C also. I dont want to give you the wrong answer though.@Vocaloid

RawlsK29:

@Vocaloid

Vocaloid:

y - y1 = m(x-x1) Let m = slope Let (x1,y1) be your point (-2,9)

joshiscool:

so B?

joshiscool:

i mean A

Vocaloid:

Think carefully

Vocaloid:

X1 = -2 y1 = 9

joshiscool:

its D?

Vocaloid:

if y 1 = 9 y - y1 =?

joshiscool:

-2

Vocaloid:

Nope Replace y1 with 9 Then re-write the expression y - y1

Vocaloid:

Leave the y alone

joshiscool:

y+9=−1/4(x−2)

Vocaloid:

Just think about my question

joshiscool:

im confused

Vocaloid:

Re expression y - y1 Erase y1 Then write Down 9 where y1 used to be Don't do anything else

Vocaloid:

*write the expression y - y1

joshiscool:

But thats not a answer option

Vocaloid:

I know that

Vocaloid:

Please follow my instructions

joshiscool:

y-9

Vocaloid:

Good So only one answer choice matches y - 9

Vocaloid:

Look and see which matches best

joshiscool:

i got a 42%

Vocaloid:

Which answer did you pick??

joshiscool:

C

Vocaloid:

Good

Vocaloid:

Redo the others if possible